- Emergency Help
- Computer Training for All Staff
- Blog Policy and Procedure
- Middle Managers Meeting Minutes
George A. Smathers Libraries
University of Florida
2124 NE Waldo Rd
Gainesville, FL 32609
Telephone: (352) 273-2800
Fax: (352) 392-2354
Information Technology Staff Meeting Notes
September 21 2010
September 21 2010
Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Winston Harris, Michael Jay, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Mark Sullivan, Evan WackAnnouncements
Review Minutes of 9/14/10
- Michael will be off for a week starting this Thursday. He is riding his motorcycle to Barber Motorsports to see an AMA superbike race in Alabama.
- Open Enrollment: 9/27/10 through 10/22/10.
- Eastside Open House: Fri, Oct 1. Announcement soon to be released.
- IT Customer Service training - October 18. Everyone is expected to attend.
Ongoing IT Updates
- Ann #3 - remove "slush".
- Non-affiliated patrons found a way around the 15-minute wait. Evan is going to create a lib-guest group and a script that will run every 2 minutes.
- Staff computer replacement - hardly any left - one in DLC that will be purchased.
- Josh will conduct training for Fiscal Services to learn how to do their own Canon meter reads.
- Win7 bug - may have figured it out. Took base and did sysprep first. It worked. Put Clean Slate in the mix, it failed. We are reimaging the computer with Clean Slate to test.
- Vaults - Laurie submitted a list to Bill Hanssen. He is out this week. When he is back, he'll submit it to Property Records. This will be the last of the IT vault items. When this load has been picked up, both vaults will be empty of IT items.
- EPO is back up (rebuilt). If staff have trouble with Site Advisor, just tell them to reboot.
- Network printer info - Josh plans on updating info in Grover and on the print server.
- Staff & public computers - deleting stale accounts in AD - Logan is working on it.
- Software tracking is very convoluted and needs to be re-hauled.
- Video conferencing - Oct 5 - a programmer will be here to work on the 1A microphones.
- SharePoint site - Hardware Unit SharePoint site demoed to IT.
- Printing - Stacey discussed with Josh the idea of getting Gator One vending card recharging stations. Unaffiliated patrons would be able to use them. We purchase, Gator One maintains.
- The Hardware Unit has worked on 370 tickets since August 23.
- Help Desk - new student, Diana Etienne started last Friday. She works MWF 1PM-5PM.
- Document disposition database - Tom and Shruti to review in-house on Friday. We will show it to Jan, etc. next week.
- Macros - there are several.
- OLE - on hold. Maureen will be returning from training next week.
- OCR automation - Kirtas documentation - Chris is trying to put together something to handle it.
- EPO is up and running.
- SharePoint - discussed in the Hardware section.
- App server move on hold. Need to come up with a move date.
- Physical server move in process. Logan went to the Library server room in East and pulled cabling from old servers. It is not quite done. NT18, 13 and two of the VM servers are still running. NT13 will stay up until GIS is installed on the new server.
- Is the switch moved? No, should be done this week.
DLC General Issues
- Baldwin ticket - Donna from FCLA provided a big text document. The Marc record helped to pull the BIB number. Turned it into an Excel spreadsheet. Cynthia needs to know to create a Marc record.
- OCR - documentation 90% finished. Connection between documents and database server tested. Turned documentation in to Laurie Taylor.
- OCR 4 - can't go away until Kirtas is fully functional. It depends on the speed. It is running at a terminal point until Kirtas is operational.
- Backups - two-tiered approach includes file server. CNS does nightly Tivoli backups, but restore is complex. We will create templates and save to shared storage. This should cover sysdrive failure. Mark will let us know if we need a new template. EPO on Altiris - if we upgrade to a new version, then we will have to create a new template. Will maintain only one template for each box.
- Link naming conventions - no request for changes.
- Oil spill site - still needs content.
- Judaica and Acquisitions - Judaica new content has not yet been received. Acquisitions should be done tomorrow.
- Proxy server - still having problems.